# Predicting the Geothermal Gradient in Colombia: a Machine Learning Approach

### v1

@@ Abstract
Accurate determination of the geothermal gradient is critical for assessing the geothermal energy potential of a given region. Of particular interest is the case of Colombia, a country with abundant geothermal resources. A history of active oil and gas exploration and production has left drilled boreholes in different geological settings, providing direct measurements of the geothermal gradient. Unfortunately, large regions of the country where geothermal resources might exist lack such measurements. Indirect geophysical measurements are costly and difficult to perform at regional scales. Computational thermal models could be constructed, but they require very detailed knowledge of the underlying geology and uniform sampling of subsurface temperatures to be well-constrained. We present an alternative approach that leverages recent advances in supervised machine learning and available direct measurements to predict the geothermal gradient in regions where only global-scale geophysical datasets and course geological knowledge are available. We find that a Gradient Boosted Regression Tree algorithm yields optimal predictions and extensively validate the trained model. We show that predictions of our model are within 12% accuracy and that independent measurements performed by other authors agree well with our model. Finnally, we present a geothermal gradient map for Colombia that highlights regions where futher exploration and data collection should be performed.
@@ 1 Highlights
  • A machine learning approach for geothermal gradient prediction using only global-scale geophysical datasets and low resolution geological maps.
@@
  • Model allows for the generation of a new geothermal gradient map of Colombia highlighting the potential of unexplored regions.
@@
  • Model predictions confirm high geothermal gradients in the Amazon, an unexplored region, giving a baseline to start studying the area.
@@ 2 Introduction
The tectonic and geological characteristics of Colombia give rise to thermal anomalies and regions exhibiting elevated geothermal gradients, which represent promising sources of geothermal energy. However, performing accurate estimations of the country’s potential for geothermal energy generation remains challenging. The most current geothermal gradient map for Colombia covers roughly half of the country’s territory (Alfaro et al., 2009), for the rest no systematic determinations of the geothermal gradient are available. Some studies have tried to address this issue. Uyeda and Watanabe (1970) observed normal to subnormal geothermal gradient values across South America, with higher values in the Andes related to geothermal activity. Bachu et al. (1995) focused on the Llanos Basin, finding that geothermal gradients decrease with depth and westward. In northwestern Colombia, Quintero et al. (2019) estimated geothermal gradients using grid systems and aeromagnetic data. Unfortunately, all of these studies are either focused on narrow areas of interest or scales that are too large for meaningful estimations of the country’s geothermal potential. Measurements from boreholes drilled by the oil and gas industry provide the best direct constraint on the geothermal gradient, in the case of Colombia they are either very sparse or clustered to specific regions, as the need to decarbonize the economy grows, it is unlikely that the past pace of oil exploration activity will continue. Therefore, a different approach is needed if Colombia is to realize its clean energy production objectives (Alfaro et al., 2021).
@@
It is seldom impossible to develop comprehensive geothermal gradient models for an entire country using only direct borehole measurements (Burton-Johnson et al., 2020), indirect approaches are necessary. Geophysical methods provide complementary estimates of the geothermal gradient over large spatial scales and depths that are inaccessible by boreholes alone; magnetic and electrical surveys exploit the thermal properties of rocks to estimate temperature gradients. While ideal, due to their cost such surveys are typically performed once a narrow target of interest has been identified. Numerical thermal modeling has emerged as a useful tool for estimating geothermal gradients as it allows for the prediction of the entire temperature field of the region’s subsurface (Békési et al., 2020). Thermal models require detailed knowledge of the underlying geology, direct temperature and heat flow measurements and involve a significant number of assumptions. For Colombia, while in principle modeling could be performed it is unlikely to yield accurate results.
@@
Data-driven approaches have become viable alternatives for the estimation of the earth’s thermal properties. Rezvanbehbahani et al. (2017) used a gradient boosting regression tree algorithm and a catalog of global geological features to produce a new geothermal heat flow (GHF) map of Greenland, their predictions were consistent with regional geology, tectonics and ice core measurements. A similar approach was used to predict GHF in Antarctica by Lösing & Ebbing (2021), it incorporated regional datasets to improve predictions. Matiz-León (2023) emphasized the importance of statistical methods and spatial prediction for estimating geothermal gradients, especially in areas lacking in-situ data. Advances in machine learning and the increasing availability of regional and global geological datasets allow for the development of cost effective approaches to provide initial baselines of the thermal properties of entire countries.
@@
In this work we present a Gradient Boosted Regression Tree algorithm to accurately predict the spatial distribution of geothermal gradients across a region. The approach only requires global scale geological and geophysical datasets, coarse geological knowledge and sparsely distributed borehole temperature measurements to produce accurate results. The paper is organized as follows: In section 2 we describe the details of the algorithm, in section 3 we present the data and the strategy used to train the model, in section 4 we provide extensive validation of our model and present a new map of geothermal gradients for Colombia. The predicted map highlights regions displaying relatively high geothermal gradients, where future exploration surveys should be conducted. Further model validation is provided by the fact that predictions show good agreement with indirect geophysical measurements of the geothermal gradient performed in the Amazon.
@@ 3 Method
Regression trees are particularly adept at identifying complex, nonlinear relationships by partitioning data recursively based on specific criteria, such as maximum depth, number of leaves, and minimum samples per leaf. In the gradient boosting framework, each iteration involves training a new tree on a subset of the data to accurately predict gradient descent steps. By aggregating these trees into an ensemble model, we enhance our capacity to predict outcomes for new data samples. This method stands out for its straightforward interpretability and operational simplicity, qualities that are often highlighted in the literature (Lösing & Ebbing, 2021).
@@ 3.1 Enhanced Gradient Boosting with XGBoost
In this study, we utilize the XGBoost algorithm, a refined version of gradient boosting regression known for its high efficiency and robust regularization techniques. Developed by Chen & Guestrin (2016), XGBoost is particularly adept at handling complex models and large datasets. The methodology involving XGBoost encompasses the following aspects:
@@
    Here, $T$ signifies the total number of leaves in tree $k$, and $w$ denotes the vector of scores on the leaves. The parameter $\gamma$ serves as the complexity cost assigned to each additional leaf, penalizing the growth of the tree structure. The term $\lambda$ is the L2 regularization parameter on the leaf weights, reducing their magnitude to enhance generalization capabilities. Collectively, these components of $\Omega(f_{k})$ ensure that the model remains both effective and interpretable.

### v7

@@ Abstract
Accurately determining the geothermal gradient is crucial for assessing geothermal energy potential. In Colombia, despite an abundance of theoretical geothermal resources, large regions of the country lack gradient measurements. This study introduces a machine learning approach to estimate the geothermal gradient in regions where only global-scale geophysical datasets and course geological knowledge are available. We find that a Gradient-Boosted Regression Tree algorithm yields optimal predictions and extensively validates the trained model, obtaining predictions of our model within 12% accuracy. Finally, we present a geothermal gradient map of Colombia that serve as an indicator of potential regions for further exploration and data collection. This map displays gradient values ranging from 16.75 to 41.20°C/km and shows significant agreement with geological indicators of geothermal activity, such as faults and thermal manifestations. Additionally, our results are consistent with independent findings from other researchers in specific regions, which supports the reliability of our approach.
@@ 1 Highlights
  • A machine learning approach for geothermal gradient prediction using only global-scale geophysical datasets and low-resolution geological maps.
@@
  • The model allows for the generation of a new geothermal gradient map of Colombia highlighting the potential of unexplored regions.
@@
  • Model predictions confirm high geothermal gradients in the Amazon, an unexplored region, giving a baseline for studying the area.
@@ 2 Introduction
The tectonic and geological characteristics of Colombia give rise to thermal anomalies and regions exhibiting elevated geothermal gradients, which represent promising sources of geothermal energy. However, performing accurate estimations of the country’s potential for geothermal energy generation remains challenging. The most current geothermal gradient map for Colombia covers roughly half of the country’s territory (Alfaro et al., 2009), for the rest no systematic determinations of the geothermal gradient are available. Some studies have tried to address this issue. Uyeda and Watanabe (1970) observed normal to subnormal geothermal gradient values across South America, with higher values in the Andes related to geothermal activity. Bachu et al. (1995) focused on the Llanos Basin, finding that geothermal gradients decrease with depth and westward. In northwestern Colombia, Quintero et al. (2019) estimated geothermal gradients using grid systems and aeromagnetic data. Matiz-León (2023) emphasized the importance of statistical methods and spatial prediction for estimating geothermal gradients in the Llanos Basin, especially in areas lacking in-situ data. Unfortunately, all of these studies are either focused on narrow areas of interest or scales that are too large for meaningful and detailed geothermal energy estimations. Measurements from boreholes drilled by the oil and gas industry provide the best direct constraint on the geothermal gradient; in the case of Colombia, they are either very sparse or clustered to specific regions, as the need to decarbonize the economy grows, it is unlikely that the past pace of oil exploration activity will continue. Therefore, a different approach is needed if Colombia is to realize its clean energy production objectives (Alfaro et al., 2021).
@@
It is seldom impossible to develop comprehensive geothermal gradient models for an entire country using only direct borehole measurements (Burton-Johnson et al., 2020), so indirect approaches are necessary. Geophysical methods provide complementary estimates of the geothermal gradient over large spatial scales and depths that are inaccessible by boreholes alone; magnetic and electrical surveys exploit the thermal properties of rocks to estimate temperature gradients. While ideal, due to their cost, such surveys are typically performed once a narrow target of interest has been identified. Numerical thermal modeling has emerged as a useful tool for estimating geothermal gradients as it allows for the prediction of the entire temperature field of the region’s subsurface (Békési et al., 2020). Thermal models require detailed knowledge of the underlying geology, direct temperature and heat flow measurements and involve a significant number of assumptions. For Colombia, while in principle modeling could be performed, it is unlikely to yield accurate results.
Data-driven approaches have become viable alternatives for the estimation of the earth’s thermal properties. Goutorbe et al. (2011) explored empirical estimators to estimate heat flow in data-scarce regions, using global geological and geophysical proxies; obtaining results that demonstrate their effectiveness in approximating heat flow. Lucazeau (2019) updates the global terrestrial heat flow dataset with around 70,000 measurements, finding that Earth’s total heat loss is approximately 40–42 TW, closely matching previous conductive cooling estimates. The study highlights improved measurement quality and coverage, especially in young oceanic regions, which significantly reduce discrepancies with existing models.
@@
Rezvanbehbahani et al. (2017) used a gradient boosting regression tree algorithm and a catalog of global geological features to produce a new geothermal heat flow (GHF) map of Greenland, their predictions were consistent with regional geology, tectonics and ice core measurements. A similar approach was used to predict GHF in Antarctica by Lösing & Ebbing (2021), it incorporated regional datasets to improve predictions. Advances in machine learning and the increasing availability of regional and global geological datasets allow for the development of cost effective approaches to provide initial baselines of the thermal properties of entire countries.
Machine learning has increasingly been used in many geothermal-related studies as well. Assouline et al. (2019) proposed a methodology to extract the very shallow geothermal potential at a national scale for Switzerland, using a combination of geographic information systems, traditional modeling, and machine learning; this approach leverages data such as monthly temperatures at various depths in the surface layer, along with thermal conductivity and diffusivity. Siler et al. (2021) applied unsupervised machine learning to identify key geological factors influencing geothermal production in Nevada’s Brady geothermal field, by analyzing geological data to discern controls on geothermal fluid pathways. Additionally, Pang et al. (2023) optimized machine learning algorithms to predict thermal conductivity (TC) using accessible, high-resolution logging data, employing techniques including random forest (RF), convolutional neural network (CNN), support vector regression (SVR), and particle swarm optimization-enhanced SVR (PSO_SVR). Furthermore, Alqahtani et al. (2023) leveraged borehole temperature and remote sensing data to pinpoint prospective zones with significant geothermal activity within the Harrat Rahat volcanic field in western Saudi Arabia, favoring exploration and drilling efforts.
@@
In this work we present a Gradient Boosted Regression Tree algorithm to accurately predict the spatial distribution of geothermal gradients across a region. The approach only requires global scale geological and geophysical datasets, coarse geological knowledge and sparsely distributed borehole temperature measurements to produce accurate results. The paper is organized as follows: In section 4 we describe the details of the algorithm, in section 5 we present the data and the strategy used to train the model, in section 6 we provide extensive validation of our model and present a new map of geothermal gradients for Colombia. The predicted map highlights regions displaying relatively high geothermal gradients, where future exploration surveys should be conducted. Further validation of the model is provided by the fact that the predictions show good agreement with indirect geophysical measurements of the geothermal gradient performed in the Amazon.
3 Geological Setting of Colombia
Colombia, located in northwestern South America, and positioned at the intersection of the Nazca, South American, and Caribbean plates, has a dynamic geological evolution extending back to 1780 million years. This dynamism is characterized by the Andean orogeny, which started in the Jurassic and is still active today, along with significant faulting and seismic activity. The Guiana Shield is part of the ancient continental core of northern South América, it consists of igneous intrusive rocks, volcanic rocks, and metamorphic rocks, such as granites, trachytes, and gneisses from the Precambrian era. The oldest rocks dated in Colombia are part of the Migmatitic Complex near San Felipe in Guainía, at the southeaast of the Eastern Llanos, with an age of 1780 million years. Similar rocks are found in several other regions in Colombia, indicating the expansive reach of the Guiana Shield across northern South America (Lobo, 1987).
The collision of the Panama Block, a segment of the Caribbean Plate, with South America in the Pliocene epoch was the event that joined the Americas. The tectonic interactions involving this block, including the collision with the South American plate, have been instrumental in shaping the Isthmus of Panama and the closure of the Central American Seaway, which had profound impacts on ocean circulation and climate. (De Porta, 2003).
The early stages of the Andean orogeny during the Jurassic and Cretaceous were marked by extensional tectonics, including rift and back-arc basin development, and the emplacement of large batholiths, associated with the subduction of cold oceanic lithosphere. In the mid to late Cretaceous (around 90 Ma), the orogenic character shifted abruptly as younger and hotter lithosphere began subducting beneath South America. This led to intense compressional deformation, Andean uplift, and subsequent erosion from the late Cretaceous onward (Ramos, 2009). The Colombian Andes are divided in three (Lobo, 1987; Gómez et al., 2020) (Figure 1):
  • The Central Cordillera (CC) is constituted by a complex geology with a diversity of rock types, igneous rocks are predominant (including the Antioquian Batholith), but metamorphic and sedimentary rocks are also present. It was formed through intense tectonic activities, including the uplifting and folding of sedimentary deposits, and is significant for its mineral wealth, particularly gold and silver. The CC features a Triassic polymetamorphic basement, intruded by plutons from the Permian to the Cenozoic, linked to the subduction of the Nazca Plate. This structure is flanked by Mesozoic intrusions and Jurassic volcaniclastic sequences, as well as exposed high-grade metamorphic rocks from previous orogenies. Cretaceous marine and volcanoclastic rocks, along with Neogene to Quaternary volcanoes, some still active, characterize the CC. The region also includes Miocene deposits and is bordered by major faults that define its tectonic limits.
  • The Eastern Cordillera (EC) is the broadest of the three and is noted for its rich deposits of coal and oil, which are the result of a sedimentary history that includes shallow marine environments from past geologic periods. Its complex geological history includes significant episodes of tectonic folding and faulting. The EC basement consists of high-grade metamorphic rocks from the Mesoproterozoic to Neoproterozoic eras, overlaid by Ordovician low-grade metamorphic rocks and Paleozoic sedimentary sequences. These are further topped with a thick succession of Cretaceous marine and Cenozoic continental sedimentary rocks, all deformed during the Andean orogeny. The region also features Jurassic sedimentary, intrusive, and volcaniclastic sequences. Its boundaries are delineated by major fault systems that mark the transition to Caguán–Putumayo and Llanos Orientales Basins (Figure 2).
  • The Western Cordillera (WC) is the lowest and shortest of the three and is marked by significant volcanic activity. This cordillera serves as a barrier between the Pacific Ocean and the rest of Colombia. It is delimited by the Cauca River valley to the east, which separates it from the Central Cordillera. The WC primarily comprises Cretaceous sedimentary rocks, gabbros, and basalts from the Caribbean-Colombian oceanic plateau, which were accreted to Colombia’s western margin from the Late Cretaceous to the Paleogene period. In its southern region, the Cordillera hosts Paleogene plutonic and volcanoclastic rocks. Towards the north, it features Miocene basalts and Pliocene volcanoclastic rocks, interspersed with minor Neogene intrusions. The range’s southern extremity contains Neogene and Quaternary volcanic deposits, with several active volcanoes that constitute the Southern Volcanic Segment of Colombia.
Figure 1: Tectonic scheme of Colombia, showing the Eastern Cordillera (EC), Central Cordillera (CC), Western Cordillera (WC); the different types of faults, the active faults, the Magdalena and Cauca rivers. Modified from Gómez et al. (2020).
CC and EC are separated by the Magdalena River Valley, which runs roughly parallel to the length of the two cordilleras and is a key feature of Colombia’s geography, hosting one of the major rivers in the country, the Magdalena River. The Middle Magdalena Valley Basin (Figure 2), part of the Magdalena River Valley, is the most extensively explored in Colombia and remains one of the most prolific areas as it hosts significant hydrocarbon resources. More than forty oil fields have been discovered within Cenozoic sediments, including Colombia’s first giant oil field, La Cira-Infantas. However, Cretaceous targets in carbonates remain underexplored. (ANH, 2010). The geological sequence in the basin begins with Jurassic continental deposits followed by Cretaceous sediments that transition from calcareous to siliciclastic, indicating a shift from transitional to marine environments. The Paleogene sequence consists predominantly of siliciclastic rocks deposited under mostly continental conditions with some marine influences. The basin has undergone three major deformational phases—rifting, thrusting, and wrenching—that have shaped various trap geometries (Barrero et al., 2007).
The Upper Magdalena Basin, which is also part of the Magdalena River Valley, is a Neogene Broken Foreland Basin, evolved from a larger Paleogene foreland basin extending eastward towards the Guyana Shield, covering an area of approximately 26,200 km2. It is flanked by Precambrian to Jurassic basement uplifts of the Eastern and Central Cordilleras. The basin’s stratigraphy begins with a Cretaceous continental sequence, followed by transgressive shales and limestones, and the Caballos sandstone unit, a key hydrocarbon reservoir. Subsequent layers include Albian to Campanian limestones, shales, and cherts, with middle Albian and Turonian organic-rich source rocks fueling all hydrocarbon accumulations. The sequence culminates with the Campanian to Maastrichtian Monserrate Formation, primarily sandstone and another significant hydrocarbon reservoir. The post-collisional Cenozoic sequence, composed of Paleogene and Neogene molasses, is entirely non-marine (Barrero et al., 2007).
The separation between the CC and the WC is known as the Cauca River Valley, it also runs parallel to the cordilleras and contains the Cauca River, which is the second major river in Colombia. It is narrower compared to the Magdalena River Valley and is flanked by volcanic peaks from the Central Cordillera. The Cauca River Valley is also noted for its gold mining, given the historical presence of gold in the surrounding geological formations, and it also has significant hydroelectric potential due to the flow of the river, supporting multiple hydroelectric plants that contribute to the energy needs of the country (Lobo, 1987).
The Llanos Basin, another significant region located in Eastern Colombia, is bounded by the Colombian-Venezuelan border to the north, the Guaicáramo fault system to the west, and the Guyana Shield to the east and the south. Its geological development began in the Paleozoic with a rifting phase, depositing siliciclastic sediments over a crystalline Precambrian basement. From the Triassic through the Late Cretaceous, it served as the eastern shoulder of a major rift system, then transitioned into a foreland basin from the Maastrichtian to the Paleocene. It has since collected thick molasse deposits up to the present (Barrero et al., 2007). The basin contains over 1700 MMBO of documented recoverable oil, including two giant fields, three major fields, and over eighty minor fields (ANH, 2010).
Figure 2: Geological map of Colombia highlighting the main basins. Modified from Gómez et al. (2020).
The Caguán-Putumayo Basin, spanning 110,304 km2 across southwestern Colombia, shares its geological history with Ecuador’s Oriente Basin as part of a foreland basin system. It is bounded by the Eastern Cordillera to the west, the Guyana Shield to the east, the Serranía de la Macarena to the north, and the international border with Ecuador to the south. The exploration of the basin has focused on structural traps involving Cretaceous and Cenozoic formations, with significant hydrocarbon prospects. The northern part of the basin, less explored, shows evidence of at least one active oil system, with structural features such as high-angle reverse faults, associated folds, and blind faults in the foothills, alongside normal faults and wedging in the foreland zone (Barrero et al., 2007; ANH, 2010).
The Vaupés-Amazonas Basin is a southeast-plunging depression located in southeastern Colombia. It comprises Lower Paleozoic rocks forming structural highs. An isolated area north of Mitú, covered by Neogene sediments, is also included within this basin. Extending south-southeast, the basin reaches the borders of Peru and Brazil, potentially connecting with the Solimoes Basin further south (Barrero et al., 2007).
3.1 Geothermal Gradient in Colombia
The scientific study of Earth’s heat has practical applications in exploiting heat as an energy source (both thermal and electric), as well as in understanding hydrocarbon and mineral resources, and modeling the evolution of Earth’s crust and tectonic processes (Alfaro et al., 2009). Colombia, located within the Pacific Ring of Fire—a tectonic belt renowned for its volcanoes and earthquakes—hosts regions that are highly favorable for generating geothermal energy. The country is home to 13 active volcanoes and numerous hot springs (Figure 3), highlighting its substantial geothermal potential (Clavijo et al., 2008; Marzolf, 2014).
Figure 3: Thermal manifestations and volcanoes in Colombia. Modified from Servicio Geológico Colombiano (SGC) (2015).
Alfaro et al. (2009) defined a preliminary geothermal gradient map to promote research on terrestrial heat flow and produce a map based on measured information. For the preliminary map, apparent geothermal gradients were calculated from temperature measurements at the maximum depth recorded (BHT) in approximately 4600 wells, primarily drilled by the oil sector, that were adjusted using the AAPG’s empirical method and included average surface temperatures. The wells, spread across 14 sedimentary basins, cover about 50% of the country’s territory (Figure 4a)
Figure 4: (a) Preliminary geothermal gradient map of Colombia. modified from Alfaro et al. (2009). (b) Colombian-Caribbean Geothermal Gradient Map. modified from Gómez et al. (2019). (c) Geothermal gradient for the Eastern-Llanos basin. modified from Matiz-Leon (2023). (c) Geothermal gradient of the Amazon basin in northern Brazil. modified from Pimentel & Hamza (2010).
More contemporary and methodologically varied studies were conducted to estimate the geothermal gradient in more localized regions, such as the one by Matiz-León (2023) for the Eastern-Llanos. To calculate the geothermal gradient, they utilized BHT measurements, applied deterministic methods such as Minimum Curvature and Inverse Weighted Distance (IDW) for initial spatial predictions, and enhanced these with probabilistic methods like Ordinary Kriging and Sequential Gaussian Simulation (SGS) for detailed 2D and 3D spatial modeling. (Figure 4c).
Gómez et al. (2019) calculated the geothermal gradient for northwestern Colombia (a region characterized by data scarcity) by first estimating the Curie point depth (CPD) using aeromagnetic data. This involved the application of spectral analysis techniques to determine the depth at which magnetic minerals lose their permanent magnetic properties due to temperature increases. The CPD was then correlated with surface heat flow measurements to derive the geothermal gradient across the region (Figure 4b).
Unfortunately, there are no studies specifically focused on calculating the geothermal gradient in the Colombian Amazon region. However, the Brazilian portion of the Amazon has been studied by Pimentel & Hamza (2010), who determined the geothermal gradients primarily using corrected Bottom Hole Temperature (BHT) data. Their study also employed Constant Bottom Hole Temperature (CBT) methods for more stable thermal fields and Conventional Logging (CVL) to calculate gradients from litological profiles (Figure 4d).
@@ 4 Method
Regression trees are particularly adept at identifying complex non-linear relationships by recursively partitioning data based on specific criteria, such as maximum depth, number of leaves, and minimum samples per leaf. In the gradient-boosting framework, each iteration involves training a new tree on a subset of the data to accurately predict gradient descent steps. By aggregating these trees into an ensemble model, we improve our ability to predict outcomes for new data samples. This method stands out for its straightforward interpretability and operational simplicity, qualities that are often highlighted in the literature (Lösing & Ebbing, 2021).
We considered several machine learning algorithms, including Random Forest Regression. Our decision to utilize the Gradient Boosted Regression Tree algorithm was primarily influenced by its superior performance in terms of error metrics when compared to alternatives.
Specifically, the eXtreme Gradient Boosted Regression Tree, or XGBoost, operates on a sequential model building process where each successive tree is trained to correct the residuals of the previous ones. This method is particularly effective in complex datasets as it methodically minimizes errors and improves prediction accuracy over iterations. In contrast, Random Forest involves training multiple decision trees independently and averaging their predictions, which can be less effective in datasets where sequential error correction is beneficial.
The enhanced performance of XGradient Boosting in handling complex patterns and interactions within our dataset justified its selection over other methods such as Random Forest, despite the latter’s robustness and popularity in various applications.
@@ 4.1 Enhanced Gradient Boosting with XGBoost
The XGBoost algorithm, used in this study, is a refined version of gradient boosting regression Developed by Chen & Guestrin (2016), known for its high efficiency and robust regularization techniques. The methodology involving XGBoost encompasses the following aspects:
@@
    Here, $T$ signifies the total number of leaves in the tree $k$, and $w$ denotes the vector of scores on the leaves. The parameter $\gamma$ serves as the complexity cost assigned to each additional leaf, penalizing the growth of the tree structure. The term $\lambda$ is the L2 regularization parameter on the leaf weights, reducing their magnitude to enhance generalization capabilities. Collectively, these components of $\Omega(f_{k})$ ensure that the model remains both effective and interpretable.